# Conefield approach to identifying regions without flux surfaces for magnetic fields

### v1

@@ 3.3 Killends
Suppose one has established a set $\mathcal{S}$ of points through which there pass no invariant torus transverse to $\xi$ and two fields of covectors $\alpha^{-}$, $\alpha^{+}$ such that the tangent plane represented by $\alpha$ to any invariant torus satisfies $\alpha^{-}\leq\alpha\leq\alpha^{+}$. To simplify the presentation, let us assume that $\alpha^{\pm}(\xi)>0$, as above.
@@
A more efficient method is to start from the boundary of $\mathcal{S}$ and enlarge $\mathcal{S}$ using the bounds on allowed slopes, as follows. For simplicity of presentation, suppose the boundary of $\mathcal{S}$ is differentiable. Take two points $x^{\pm}$ of $\partial\mathcal{S}$ connected by an arc of $\partial\mathcal{S}$ such that $\mathrm{ker}\,\alpha^{+}$ is tangent to $\partial\mathcal{S}$ at $x^{+}$ and $\mathrm{ker}\,\alpha^{-}$ is tangent to $\partial\mathcal{S}$ at $x^{-}$, see figure 3. Integrate $\mathrm{ker}\,\alpha^{\pm}$ leftwards in the chosen section until they intersect. The region they bound with $\partial\mathcal{S}$ can be added to $\mathcal{S}$. The similar procedure applies at the right of $\mathcal{S}$. In the present work, we will refer to the curves created by the integration of $\mathrm{ker}\,\alpha^{\pm}$ as ‘killends branches’.
@@ 4 Results
It is important to highlight that areas are equivalent whether computed in symplectic coordinates $(\tilde{y},\tilde{z})$ or in $(\psi,\vartheta)$, since $d\tilde{y}\wedge d\tilde{z}=d\psi\wedge d\vartheta$. Consequently, the area $S$ of the non-existence region in the plane $\phi=0$ can be estimated by counting the number of initial conditions detected by the method on a regular grid in $(\tilde{y},\tilde{z},\phi=0)$. In other words, if $\mathcal{S}$ represents the set of points identified by the conefield formulation on an $N\times N$ regular grid over the domain $[\tilde{y}_{0}-L,\tilde{y}_{0}+L]\times[\tilde{z}_{0}-L,\tilde{z}_{0}+L]$, the area $S$ is approximated by
@@
For a given trajectory, the interval of integration is reported as $t_{*}$ if the time $t^{*}$ at which non-existence was detected, or $t_{f}$ if it was not detected. As a measure of non-existence of tori of a given class, figures display the relative time of detection shown in hues, using the ratio $q=t_{*}/t_{f}$, which represents the time of detection relative to the timeout.
In all the examples throughout this section and elsewhere, we take the following values and function for the vector potential (4)
@@
As previously mentioned, the results in all forthcoming figures are presented over the poloidal plane $\phi=0$.
@@ 4.1 Conefield for a non-integrable example
Figure 4: Conefield results (shown in hues) and three zoomed-in insets (on the left) displaying the computed conefield on the complement of the detected non-existence region $\tilde{\mathcal{S}}$, on a regular grid of initial conditions in the positive quadrant of the poloidal plane $\phi=0$ for the magnetic field derived from (13) with $\varepsilon=0.003$ in symplectic coordinates. The hues vary from fast detection (red) to no detection at all (blue) within the timeout $t_{f}=80$. $q$ denotes the ratio of the time to detection to the timeout, except that $q=1$ denotes no detection within the timeout.
@@
Figure 7 shows the computed area $S=S(t_{f})$ of nonexistence from the conefield approach compared with the results in [5] for the same example for two different values of $\varepsilon$. It can be seen that the two implementations yield similar results. In theory, we would expect the results to yield almost identical plots for the same integration time. The difference may imply that the initialization values are not equivalent for both formulations. Regardless, the results are similar for a modest integration interval of $t_{f}=300\sim 48$ laps around the $z$-axis.
Figure 7: Nonexistence area $S(t_{f})$ detected for two different values of $\varepsilon$ for the field (13), by two Converse KAM methods: conefield approach (solid line) and direct approach in [5].
@@ 4.2 Killends in a non-integrable example
Figure 8: Killends continuation branches and area regions for the upper plane of the example in symplectic coordinates $(\tilde{y},\tilde{z})$ for the perturbation parameter value $\varepsilon=0.002$ and $t_{f}=100$.
@@
Indeed, in cases with positive shear, no fieldline of minimal action can be eliminated by the direct or conefield methods, and this includes not only those on invariant tori of the given class, but also the minimising hyperbolic periodic orbits and the fieldlines on minimising cantori, but killends can eliminate minimising orbits if they do not lie on invariant tori. Thus, the killends extension to the conefield method makes an important contribution by eliminating more. Even though the area of the extra eliminated points might be negligible, the time taken to eliminate them is greatly reduced.
@@ 5 Conclusions
In conclusion, we have reformulated the Converse KAM method applied to magnetic fields in [11], in terms of construction of a conefield, and extended the non-existence region using the supplementary “killends” condition. We find that using the killends extension we can eliminate more volume. Furthermore, we find that we can eliminate more volume with less computation time.
As in [5], the direction field used in the computations was radial, $\xi=\partial_{\psi}$, to eliminate points not in the principal class. However, it is possible to replace it with a vector field that allows us to identify points in the magnetic island.
Each method has its trade-offs. By comparing the results from the conefield approach to the direct approach in [5], we can verify the accuracy of our output to increase confidence in their application, and evaluate strengths and weaknesses in each method. Even though both methods have comparable efficiency in computation steps and running time, the conefield approach yields additional information on the non-eliminated points, which is key to allow us to construct the killends branches and hence to eliminate more points.
The killends branches computation presented in this paper, used the general knowledge of the structures of the magnetic islands chosen. In a more general case, in which the location or symmetry of the islands is not known beforehand, the implementation would require additional strategies to select suitable seed points.
However, the killends results show that the upper bounds computed for the non-existence area saturate already for fairly short timeout, which translates to close to optimal bounds for low timeouts.
@@ Acknowledgements
We are grateful to Nikos Kallinikos for comments. This work was supported by a grant from the Simons Foundation (601970, RSM).

### v2

@@ 3.3 Killends
Suppose one has established a set $\mathcal{S}$ of points through which there pass no invariant torus transverse to $\xi$ and two fields of covectors $\alpha^{-}$, $\alpha^{+}$ on the complement of $\mathcal{S}$ such that the tangent plane represented by $\alpha$ to any invariant torus satisfies $\alpha^{-}\preceq\alpha\preceq\alpha^{+}$. To simplify the presentation, let us assume that $\alpha^{\pm}(\xi)>0$, as above.
@@
A more efficient method is to start from the boundary of $\mathcal{S}$ and enlarge $\mathcal{S}$ using the bounds on allowed slopes, as follows. For simplicity of presentation, suppose the boundary of $\mathcal{S}$ is differentiable. Take two points $x_{1},x_{2}$ of $\partial\mathcal{S}$ connected by an arc of $\partial\mathcal{S}$ such that $\mathrm{ker}\,\alpha^{+}$ is tangent to $\partial\mathcal{S}$ at $x_{1}$ and $\mathrm{ker}\,\alpha^{-}$ is tangent to $\partial\mathcal{S}$ at $x_{2}$, see figure 4. Integrate $\mathrm{ker}\,\alpha^{\pm}$ leftwards in the chosen section until they intersect. The region they bound with $\partial\mathcal{S}$ can be added to $\mathcal{S}$. The similar procedure applies at the right of $\mathcal{S}$. In the present work, we will refer to the curves created by the integration of $\mathrm{ker}\,\alpha^{\pm}$ as ‘killends branches’.
@@ 4 Results
It is important to highlight that areas are equivalent whether computed in symplectic coordinates $(\tilde{y},\tilde{z})$ or in $(\psi,\vartheta)$, since $d\tilde{y}\wedge d\tilde{z}=B_{0}^{-1}d\psi\wedge d\vartheta$. Consequently, the area $S$ of the non-existence region in the plane $\phi=0$ can be estimated by counting the number of initial conditions detected by the method on a regular grid in $(\tilde{y},\tilde{z},\phi=0)$. In other words, if $\mathcal{S}$ represents the set of points identified by the conefield formulation on an $N\times N$ regular grid over the domain $[\tilde{y}_{0}-L,\tilde{y}_{0}+L]\times[\tilde{z}_{0}-L,\tilde{z}_{0}+L]$, the area $S$ is approximated by
@@
For a given trajectory, the interval of integration is reported as the time $t_{*}$ at which non-existence was detected, or $t_{f}$ if it was not detected. As a measure of non-existence of tori of a given class, figures display the relative time of detection shown in hues, using the ratio $q=t_{*}/t_{f}$, which represents the time of detection relative to the timeout.
In all the examples throughout this section and elsewhere, we take the following values and functions for the vector potential (4)
@@
We select only finitely many of the $\varepsilon_{mn}$ to be non-zero. As previously mentioned, the results in all forthcoming figures are presented over the poloidal plane $\phi=0$.
@@ 4.1 Conefield for a non-integrable example
Figure 5: Conefield results (shown in hues) and three zoomed-in insets (on the left) displaying the computed conefield on the complement of the detected non-existence region $\tilde{\mathcal{S}}$, on a regular grid of initial conditions in the positive quadrant of the poloidal plane $\phi=0$ for the magnetic field derived from (14) with $\varepsilon=0.003$ in symplectic coordinates. The hues vary from fast detection (red) to no detection at all (blue) within the timeout $t_{f}=80$ (which corresponds to about 15 toroidal revolutions). $q$ denotes the ratio of the time to detection to the timeout, except that $q=1$ denotes no detection within the timeout.
@@
Figure 8 shows the computed area $S=S(t_{f})$ of nonexistence from the conefield approach compared with the results in [5] for the same example for two different values of $\varepsilon$. It can be seen that the two implementations yield similar results. In theory (compare [3]), we would expect the results to yield almost identical plots for the same integration time, because the methods are equivalent except for the time-interval of integration: $[0,t_{f}]$ for [5] versus $[-t_{f}/2,t_{f}/2]$ for the conefield method. But as our initial conditions are only on $\phi=0$, this does mean that a slightly different sample of phase space is selected. The difference in results may reflect this difference in the origin of the time-interval. Regardless, the results are similar for a modest integration interval of $t_{f}=300\sim 48$ laps around the $z$-axis.
Figure 8: Nonexistence area $S(t_{f})$ detected for three different values of $\varepsilon$ for the field (14), by two Converse KAM methods: conefield approach (solid line) and direct approach in [5].
@@ 4.2 Killends in a non-integrable example
Figure 9: Killends continuation branches and area regions for the upper plane of the example in symplectic coordinates $(\tilde{y},\tilde{z})$ for the perturbation parameter value $\varepsilon=0.002$ and $t_{f}=100$. The enclosed area is displayed in light yellow.
@@
[…]
5 Discussion
The magnetic fields we have studied here have a reflection symmetry (“stellarator symmetry”) [11]. It could be used to speed up computations for initial conditions on the symmetry lines by a factor of two, as was done in [3] and [5]. We didn’t implement it here, to avoid obscuring the general procedure.
The killends branches computation presented in this paper, used some prior knowledge of the structures of the magnetic islands chosen. For a more general case, in which the location or symmetry of the islands is not known beforehand, the implementation would require additional strategies to select suitable seed points. A challenge for the future is to automate the killends procedure. This was achieved for area-preserving maps in [3] by subdividing the domain into squares and computing upper and lower bounds on the slopes in each square; then killends was applied in an automatic way to eliminate whole squares where possible. Something similar could be done here but it would be preferable not to take bounds on a subdivision of the domain.
To simplify our presentation of the conefield method, we considered only differentiable tori of the given class. But actually, the method applies equally well to all Lipschitz tori topologically transverse to the given direction field. A torus is Lipschitz if it is locally the graph $z=f(x,y)$ of a Lipschitz function $f$, i.e. for any point $(x_{0},y_{0},z_{0})$ of the torus there is a neighbourhood in which the intersection with the torus is contained within $|z-z_{0}|\leq L(|x-x_{0}|+|y-y_{0}|)$ for some $L$. Then $\alpha^{\pm}$ constructed by the conefield method provide local Lipschitz bounds on the intersection of an invariant torus with a transverse section through the given point. This is a particularly useful result in case of definite shear in a domain of the form of the product of a torus $\mathcal{T}$ with an interval, with direction field $\xi$ in the interval direction, because translating a theorem of Birkhoff (sec.44 of [12]) to this context shows that every invariant torus that is homotopic to $\mathcal{T}\times{c}$, $c$ a constant, is transverse to $\xi$ and Lipschitz. So the class of tori that the method treats can be considered to be all those homotopic to $\mathcal{T}\times{c}$.
As in [5], the direction field used in the computations was radial, $\xi=\partial_{\psi}$. This was chosen to eliminate points not on tori in the “principal” class that surround the magnetic axis, on the assumption that all such tori are transverse to the radial field. As seen in the previous paragraph, this is justified in the case of definite shear. Furthermore, changing the direction field a little does not change the definiteness of the shear, so does not change the class of tori eliminated. But the assumption might not be justified in general. For example, if the shear changes sign in the domain then there might be “wiggly” tori as found in [13].
Moreover, it is possible to replace the radial direction field with a vector field that allows tori in one or both of the magnetic islands in our examples too, as was done for an analogous problem in [14]. Our aim here was just to demonstrate use of the conefield method and killends extension in some simple cases; we leave more interesting applications for future work.
An important direction for future work is to apply the method to guiding-centre motion in a magnetic field. This consists of a two-parameter family of 3D volume-preserving vector fields, with parameters the energy and the magnetic moment. The direction field needs choosing to correspond to given class of guiding-centre motion. For volume-form on non-degenerate energy levels one can take $\frac{e\tilde{B}_{\parallel}}{v_{\parallel}}\Omega$, with notation as in [15] (despite appearances this is non-singular).
@@ 6 Conclusions
In conclusion, we have reformulated the Converse KAM method that was applied to magnetic fields in [5], in terms of construction of a conefield, and extended the non-existence region using the supplementary “killends” condition. We find that using the killends extension we can eliminate more volume. Furthermore, we find that we can eliminate more volume with less computation time.
The conefield method is slightly more complicated to implement than the direct method of [5] but has the advantage of producing more information, namely the conefield, giving bounds on the slopes of all invariant tori of the given class and permitting the killends extension.
The killends extension allows to eliminate many points more quickly and to eliminate some other points that would never be eliminated without it. If implemented optimally, it is plausible that killends gets arbitrarily close to the true non-existence region as the timeout is increased (compare [10]). The results of our paper indicate that the non-existence region using killends saturates already for fairly short timeout in reasonable examples.
@@ Acknowledgements
We are grateful to Nikos Kallinikos for comments and to the reviewers for their careful reading, questions and suggestions. This work was supported by a grant from the Simons Foundation (601970, RSM).